# Neil A. Chriss

### v2

@@ 6.2 Naive centralization implementation costs
Continuing with the same setup as above, we will explicitly compute the cost of each of the four strategies above, that is for the aggregated firm and non-firm strategies for the cooperative and non-cooperative scenarios in Figures 7 and8.
We will compare the costs of each approach. First compute the non-cooperative firm implementation cost, specifically this covers the case when traders $1,\dots,n_{1}$ trade independently of one another. This is nothing more than the sum of the costs in (5.4):
Firm cost ($\begin{split}\textit{Firm cost (non-cooperative)}&=\sum_{i=1}^{n_{1}}\text{Cost}(\lambda_{i}a_{i},m)\\ &=\sum_{n=1}^{n_{1}}\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}\\ &=\kappa\frac{\lambda_{n_{1}}n-n_{1}}{n(1-e^{-\kappa})}+\alpha\frac{n_{1}}{n(e^{\alpha}-1)}+\frac{n_{1}\kappa}{n+1}\end{split}$ (6.6)
where as usual $\alpha=\kappa\frac{n-1}{n+1}$. We can similarly compute the non-firm traders’ cost in the non-cooperative case:
Non-firm cost ($\begin{split}\textit{Non-firm cost (non-cooperative)}&=\sum_{i=n_{1}+1}^{n}\text{Cost}(\lambda_{i}a_{i},m)\\ &=\sum_{n=n_{1}+1}^{n}\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}\\ &=\kappa\frac{\lambda_{n_{2}}n-n_{2}}{n(1-e^{-\kappa})}+\alpha\frac{n_{2}}{n(e^{\alpha}-1)}+\frac{n_{2}\kappa}{n+1}\end{split}$ (6.7)
where $\lambda_{n_{2}}=\sum_{i=n_{1}+1}^{n}\lambda_{i}$ represents the fraction of the total target quantity sought be the non-firm traders. Now we can compute the same quantities in the cooperative case. On the other hand, when the firm trades are centralized, the firm trades the cooperative firm strategy in which case the total number of traders is $n_{2}+1$ consisting of the $n_{2}$ non-firm traders and the firm. The firm’s strategy is labeled $a_{n_{1}}$. The cost of implementing the strategy $\lambda_{n_{1}}a_{n_{1}}$ is calculated applying Eq. 5.4 setting the number of traders to $n_{2}+2$ and replacing $\alpha$ with $\hat{\alpha}=\kappa\frac{n_{2}}{n_{2}+2}$:
@@
Non-firm costs ($\begin{split}\textit{Non-firm costs (cooperative)}&=\sum_{i=n_{1}+1}^{n}\kappa\frac{\lambda_{i}(n_{2}+1)-1}{(n_{2}+1)(1-e^{-\kappa})}+\hat{\alpha}\frac{1}{(n_{2}+1)(e^{\hat{\alpha}}-1)}+\frac{\kappa}{n_{2}+2}\\ &=\kappa\frac{\lambda_{n_{2}}(n_{2}+1)-n_{2}}{(n_{2}+1)(1-e^{-\kappa})}+\hat{\alpha}\frac{n_{2}}{(n_{2}+1)(e^{\hat{\alpha}}-1)}+\frac{\kappa n_{2}}{n_{2}+2}\end{split}$ (6.9)
@@ 6.3 Detailed numerical examples of naive centralization
We now provide a detailed examination of the difference in cost when a firm centralizes its trades based on the implementation cost formulas presented above. As in the prior section there is a firm with $n_{1}$ traders, $n_{2}$ other traders and (the non-firm traders) and the total target quantity is scaled to one. The firm’s portion of the target quantity is $\lambda_{n_{1}}$. This section provides data and analysis of the change in implementation costs to the firm and the non-firm traders in the aggregate, when the firm switches to centralized trading. Note that the general picture may seem counter-intuitive, as firm’s do not always or generally experience lower implementation costs (at least as far as the models go) when they centralize, because as we will see the firm becomes a target for front-running as its trades become larger and more predictable (see Section 2.3). Note also that in Section 6.4 we look at at related and more general scenarios where a firm disguises the number of traders it actually has in order to reduce implementation costs.
@@
The big picture may be summarized as when a firm centralizes its trades the total implementation over all traders decreases but the benefits accrue to the non-firm traders and not the firm. As this is undoubtedly a surprising result it’s important to pin down precisely what the underlying assumptions are. For starters, for this to hold the market must realize that the firm is trading as a single monolithic block and adjust to this to form a new equilibrium. With this said, we make the following comments regarding this. When a firm centralizes its trades the following occur:
@@
$\lambda_{n_{1}}$ $n_{1}$ $\kappa$ Firm Non-firm Total Non-Coop Coop Non-Coop Coop
@@
Table 2: The average value of change in implementation costs for a variety of scenarios when a firm switches from non-cooperative trading to centralized trading. In each case the total number of traders is one of 20, 21 and 22 and the average number of traders in the firm in this table is always 4, meaning the traders represent a minority of all traders. Each block of rows is a variety of scenarios for a fixed $\kappa$ and what varies is the fraction of the total target quantity being that the firm represents, denoted as $\lambda_{1}$. The percent changes columns display the change in implementation cost switching for the firm, for the aggregate of the non-firm and firm traders.
@@
$\lambda_{n_{1}}$ $n_{1}$ $\kappa$ Firm Non-firm Total Non-Coop Coop Non-Coop Coop
@@
Table 3: The average value of change in implementation costs for a variety of scenarios when a firm switches from non-cooperative trading to centralized trading. In each case the total number of traders is one of 20, 21 and 22 and the average number of traders in the firm in this table is always 15, meaning the traders represent a majority of all traders. Each block of rows is a variety of scenarios for a fixed $\kappa$ and what varies is the fraction of the total target quantity being that the firm represents, denoted as $\lambda_{1}$.The percent changes columns display the change in implementation cost switching for the firm, for the aggregate of Independents and in the aggregate.
@@ 6.4 Strategic centralization
Given the observations above, namely that centralizing trades counter-intuitively increases implementation costs, it stands to reason that reversing the process, namely of de-centralizing trades in some manner might decrease costs. Can a firm even go further and disguise the true number of traders it has to fool the market into believing it has more than it does. We call this process disguising order flow and examine in what circumstances and to what extent it can improve implementation costs.
examine the change in implementation costs when a firm parcels out its trades in order to appear as if the number of non-firm traders in the market is different than it actually is. We assume we are in the same situation as when studying trade centralization (see Section 6). There is a firm with $n_{1}$ traders whose target quantity in the aggregate is $\lambda_{n_{1}}$ and there are $n_{2}$ traders outside of the firm (the non-firm traders) whose target quantities in the aggregate are $\lambda_{n_{2}}$. The total number of traders is $n=n_{1}+n_{2}$.
@@
Now imagine that the wants to "pretend" that the firm sends its orders out in such a way as change the number of traders from $n_{1}$ to $n_{1}+m$ where $m+n_{1}\geq 1$ so that the firm masks the real number of traders it has and makes it appear as if there are $n_{1}+m$ so that if $m<0$ it is reducing the number of trades and if $m>0$ it is increasing the number of trades. In all cases the firm still has a total target quantity of $\lambda_{n_{1}}$. Thus the firm has disguised its true number of traders by altering the umber of traders. In this case we have:
@@
and therefore firm aggregate implementation cost when the number of traders goes from $n_{1}$ to $n_{1}+m$:
@@
Note that in theory we can take the derivative of Eq. 6.12 with respect $m$ and calculate where its zeros are in order to see at what value of $m$ costs are minimized, however we will only explore this numerically. One thing that we can do, however, is let $m\to\infty$ we see that $\hat{\alpha}\to\kappa$ and $(n_{1}+m)/(n+m)\to 1$ therefore the first term in Eq. 6.10 tends to $\frac{\kappa\lambda_{n_{1}}-\kappa}{1-e^{-\kappa}}$, while the second and third terms tend to $\frac{\kappa}{e^{\kappa}-1}$ and $\kappa$ respectively. Therefore as $m\to\infty$ the firm implementation cost tends to a limiting value of:
We can perform a similar procedure for the non-firm trader’s implementation costs where $n_{2}$ is the number of traders not in the firm. Here the situation is particularly simple because the $n_{2}/n\to 0$ as the number of splits grows large. Therefore the non-firm traders’ costs tends to
[…]
and therefore in the limit as the number of traders grows large we have the non-firm traders’ (non-firm traders) aggregate implementation costs tends to $\frac{\kappa\lambda_{n_{2}}}{1-e^{-\kappa}}$ as the number of traders grows large,
@@ 6.5 Visualizations of strategic centralization implementation costs
The firm has one trader and is a large fraction of all traders: we make three plots (each with three sub-plots) in which the firm always has one actual trader ($n_{1}=1$) and the total number of traders is 2, 3 and 10 so that the firm is 1/2, 1/3 and 1/10th of the total number of traders. We then examine implementation cost changes for three levels of alpha-decay, $\kappa=1,2.5$ and $5$. There are a few obviously interesting trends.
Firm is most of the traders: Below we see the case where the firm can substantially benefit from reverse disguising and making it appear as if there are few traders than the actually are. In these plots the black dot represents the baseline in which the firm has not tried to disguise the number of traders.
The firm has one trader in a crowded trade: In the next set of plots the firm has only one trader. The question is whether or not the firm can benefit from disguising its order flow to lower costs. The answer has similar shape to the examples above. As in many, all of the gains occur when adding only a few traders and then very rapidly the gains dissipate and ultimately costs increase.
The firm represents two-thirds of all traders: in the plots below the firm’s total number of actual traders varies: $n_{1}=10,20,50$ from left to right respectively and in all cases the total number of traders is 3/2 of the firm’s so that the firm is 2/3rds of the total. In these cases adding traders generally raises implementation costs and the firm benefits from making it appear as if it has fewer traders than it actually does. In the plots below, negative values in the x-axis represent when the firm has consolidated trades and the black dot is the baseline when the firm has not disguised its number of traders

### v3

@@ 6.2 Naive centralization implementation costs
Continuing with the same setup as above, we will explicitly compute the cost of each of the four strategies above, that is for the aggregated firm and non-firm strategies for the "with centralization" and "no centralization" cases in Figures 6 and7.
We will compare the costs of each approach. First compute the implementation costs in the case where there is no centralization, specifically this means traders $1,\dots,n_{1}$ trade independently of one another. This is nothing more than the sum of the costs in (5.4):
Firm costs (no $\begin{split}\textit{Firm costs (no centralization)}&=\sum_{i=1}^{n_{1}}\text{Cost}(\lambda_{i}a_{i},m)\\ &=\sum_{i=1}^{n_{1}}\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}\\ &=\kappa\frac{\lambda_{n_{1}}n-n_{1}}{n(1-e^{-\kappa})}+\alpha\frac{n_{1}}{n(e^{\alpha}-1)}+\frac{n_{1}\kappa}{n+1}\end{split}$ (6.6)
where as usual $\alpha=\kappa\frac{n-1}{n+1}$. We can similarly compute the non-firm traders’ cost in the no-centralization case:
Non-firm costs (no $\begin{split}\textit{Non-firm costs (no centralization)}&=\sum_{i=n_{1}+1}^{n}\text{Cost}(\lambda_{i}a_{i},m)\\ &=\sum_{i=n_{1}+1}^{n}\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}\\ &=\kappa\frac{\lambda_{n_{2}}n-n_{2}}{n(1-e^{-\kappa})}+\alpha\frac{n_{2}}{n(e^{\alpha}-1)}+\frac{n_{2}\kappa}{n+1}\end{split}$ (6.7)
where $\lambda_{n_{2}}=\sum_{i=n_{1}+1}^{n}\lambda_{i}$ represents the fraction of the total target quantity sought be the non-firm traders. Now we can compute the same quantities in the cooperative case. On the other hand, when the firm trades are centralized, the market-wide number of traders is $n_{2}+1$ consisting of the $n_{2}$ non-firm traders and one "consolidated" firm trader. The firm’s strategy is labeled $a_{n_{1}}$. The cost of implementing the strategy $\lambda_{n_{1}}a_{n_{1}}$ is calculated using Eq. 5.4 setting the number of traders to $n_{2}+2$ and replacing $\alpha$ with $\hat{\alpha}=\kappa\frac{n_{2}}{n_{2}+2}$:
@@
Non-firm costs (w/ $\begin{split}\textit{Non-firm costs (w/ centralization)}&=\sum_{i=n_{1}+1}^{n}\kappa\frac{\lambda_{i}(n_{2}+1)-1}{(n_{2}+1)(1-e^{-\kappa})}+\hat{\alpha}\frac{1}{(n_{2}+1)(e^{\hat{\alpha}}-1)}+\frac{\kappa}{n_{2}+2}\\ &=\kappa\frac{\lambda_{n_{2}}(n_{2}+1)-n_{2}}{(n_{2}+1)(1-e^{-\kappa})}+\hat{\alpha}\frac{n_{2}}{(n_{2}+1)(e^{\hat{\alpha}}-1)}+\frac{\kappa n_{2}}{n_{2}+2}\end{split}$ (6.9)
@@ 6.3 Detailed numerical examples of naive centralization
We now provide a detailed examination of the difference in cost when a firm centralizes its trades based on the implementation cost formulas presented above. As in the prior section there is a firm with $n_{1}$ traders, $n_{2}$ other traders and (the non-firm traders) and the total target quantity is scaled to one. The firm’s portion of the target quantity is $\lambda_{n_{1}}$. This section provides data and analysis of the change in implementation costs to the firm and the non-firm traders in the aggregate, when the firm switches to centralized trading. Note that the general picture may seem counter-intuitive, as firm’s do not always or generally experience lower implementation costs (at least as far as the models go) when they centralize, because as we will see the firm becomes a target for front-running as its trades become larger and more predictable (see Section 2.3). In Section 6.4 we look at related and more general scenarios where a firm masks the number of traders it actually has, thus representing a different number to the market, with the express intention of reducing its implementation costs.
@@
The big picture may be summarized as when a firm centralizes its trades the total implementation over all traders decreases but the benefits accrue to the non-firm traders and not the firm. As this is undoubtedly a surprising result it’s important to pin down precisely what the underlying assumptions are. For starters, for this to hold the market must realize that the firm is trading as a single monolithic block and adjust to this to form a new equilibrium. With this said, we make the following comments regarding this. When a firm naivelycentralizes its trades the following happens with respect to implementation costs:
@@
$\lambda_{n_{1}}$ $n_{1}$ $\kappa$ Firm Non-firm Total No Central. W/ Central. No Central. W/ Central.
@@
Table 2: The average value of change in implementation costs for a variety of scenarios when a firm switches from no centralization to centralized trading. In each case the total number of traders is one of 20, 21 and 22 and the average number of traders in the firm in this table is always 4, meaning the traders represent a minority of all traders. Each block of rows is a variety of scenarios for a fixed $\kappa$ and what varies is the fraction of the total target quantity being that the firm represents, denoted as $\lambda_{1}$. The percent changes columns display the change in implementation cost switching for the firm, for the aggregate of the non-firm and firm traders.
@@
$\lambda_{n_{1}}$ $n_{1}$ $\kappa$ Firm Non-firm Total No Central. W/ Central. No Central. W/ Central.
@@
Table 3: The average value of change in implementation costs for a variety of scenarios when a firm switches from no centralized trading to centralized trading. In each case the total number of traders is one of 20, 21 and 22 and the average number of traders in the firm in this table is always 15, meaning the traders represent a majority of all traders. Each block of rows is a variety of scenarios for a fixed $\kappa$ and what varies is the fraction of the total target quantity being that the firm represents, denoted as $\lambda_{1}$.The percent changes columns display the change in implementation cost switching for the firm, for the aggregate of Independents and in the aggregate.
@@ 6.4 Strategic centralization
Given the observations above, namely that centralizing trades counter-intuitively increases firm implementation costs, it stands to reason that reversing the process, namely of splitting trades up in some manner might decrease firm implementation costs. Put differently, can a firm even go further and mask the true number of traders so as to deceive the market into believing there are more traders competing than there actually are. We call this process misrepresenting order flow and examine in what circumstances and to what extent it can improve implementation costs. We examine the change in implementation costs when a centralizes its trades and re-arranges them to make the market believe the number traders competing is different than it actually is.
Before proceeding a small example is in order to illustrate the key points. suppose there are four traders in the firm and eight non-firm traders for a total of 12 traders taking up the trade. Then The firm has the option to naively centralize its four traders making the apparent total number of traders taking up the trade to be nine (one from the firm and the eight non-firm traders). On the other hand, the firm, after centralizing, can split the total trades into as many trades as it wants and either route to independent brokers unaware of the total, or trade as if they are independent. As an example, the firm could split the consolidated order into twenty independent (and much smaller) trades, each 5% of the firm’s total target quantity. There are two important points in this:
  • When the firm centralizes its trade to modify represent a different number of traders than it actually has, what matters is that the total number of traders competing for the stock is modified as well. This changes the cost over all traders if trading. If the firm "adds" traders by representing more than it actually has, then total cost goes up, if it "reduces" traders, then cost goes down; and
  • Through all of this, the firm never changes its target quantity. The firm’s target quantity is taken to be exogenously given by the aggregate of its traders.
We assume we are in the same situation as when studying trade centralization (see Section 6). There is a firm with $n_{1}$ traders whose total target quantity over all traders is $\lambda_{n_{1}}$. There are $n_{2}$ traders outside of the firm (the non-firm traders) whose target quantities in the aggregate are $\lambda_{n_{2}}$. The total number of traders is $n=n_{1}+n_{2}$. Note that the cost to the firm with no centralization is given by:
@@
Now imagine the firm centralizes its trades and divides its orders into $n_{1}+m$ separate parent orders, where $n_{1}+m\geq 1$. That is, the firm uses the fact that its traders are not totally autonomous to mask the true number of traders it has. This in effect is changing the game that all traders are playing. If $m<0$ the firm is reducing the apparent number of traders and if $m>0$ it is increasing the number of traders. Note that the firm will do this without modifying the total quantity it seeks to trade.
None of this is to say that non-firm traders know (or care) how many traders the firm has. In order to trade in equilibrium each trader needs to know the total number of traders taking up the trade and what their fractional target quantity is. Thus the firm’s aim is to mask its true number of traders so as to change the total number of traders taking up the trade. In this case we have:
@@
and therefore the firm’s aggregate implementation cost when centralizing trades and modifying by $m$ traders is given by (after a little rearranging of Eq. 6.10):
$\kappa\frac{\lambda_{n_{1}}(n+m)-(n_{1}+m)}{(n+m)(1-e^{-\kappa})}+\kappa\frac{n_{1}+m}{n+m+1}+\hat{\alpha}\frac{n_{1}+m}{(n+m)(e^{\hat{\alpha}}-1)}$ (6.12)
With this in hand we can provide provide insight into what a firm’s optimal behavior should be concerning centralization.
Optimal strategic centralization: The question is, for what $m$ is the implementation cost in Eq. 6.12 least? If one notes that for relatively small values of total traders $n+m$ we have $\hat{\alpha}\approx\kappa$ then we can approximate (6.12) by
$\kappa\frac{\lambda_{n_{1}}(n+m)-(n_{1}+m)}{(n+m)(1-e^{-\kappa})}+\kappa\frac{n_{1}+m}{(n+m)(e^{\kappa}-1)}+\kappa\frac{n_{1}+m}{n+m+1}$ (6.13)
and then it is straightforward to show that Eq. 6.13 the arg-min of Eq. 6.13 is achieved at:
$m=-n_{1}+\sqrt{n_{2}(n_{2}+1)}$ (6.14)
Put differently, the optimal number of traders for a firm to represent to the market in order to minimize implementation cost is given by $n_{1}+m$:
$m+n_{1}=\sqrt{n_{2}(n_{2}+1)}$ (6.15)
We prove this in Appendix B. In other words, to optimally reduce implementation costs a firm needs to centralize its trades and then modify its total number of traders by adding $m$ traders. Thus we conclude in most cases the optimal number of traders for the firm to represent when centralizing is between $n_{2}$ and $n_{2}+1$. We note a few important aspects of Eq. 6.14:
  • Balanced number of traders: strategic centralization minimizes implementation cost when the number of traders inside the firm is set to approximately the number outside. This is because the modification number $m\approx-n_{1}+n_{2}$, therefore the new number of traders in the firm is $n_{1}+m\approx n_{2}$ (one can see this visually in Figures 8 to 11;
  • Target quantity and alpha-decay do not matter: at least to first order, the number of trader’s for the firm to represent in order to minimize implementation costs is independent of both the firm’s aggregate target quantity (over all its traders) and the alpha-decay parameter $\kappa$. This is clearly the case because Eq. 6.14 simply has no terms involving $\lambda_{n_{1}}$ or $\kappa$, however we can also see this visually in Figures 8 to 11. Finally in Appendix B one can see specifically in the derivation of the proof of Eq. 6.14.
It’s worth pointing out that it is somewhat surprising that the optimal number of tradings to end up with in strategic centralization is the same number of non-firm traders and that this is independent of the firm’s target quantity and alpha-decay. It is also worth asking whether or not this is an artifact of features of the model itself such as the transaction cost models.
Limiting quantities as splitting grows: One thing that we can do, however, is let $m\to\infty$ we see that $\hat{\alpha}\to\kappa$ and $(n_{1}+m)/(n+m)\to 1$ therefore the first term in Eq. 6.10 tends to $\frac{\kappa\lambda_{n_{1}}-\kappa}{1-e^{-\kappa}}$, while the second and third terms tend to $\frac{\kappa}{e^{\kappa}-1}$ and $\kappa$ respectively. Therefore as $m\to\infty$ the firm implementation cost tends to a limiting value of:
We can perform a similar procedure for the non-firm trader’s implementation costs where $n_{2}$ is the number of traders not in the firm. Here the situation is particularly simple because as the number of traders added grows, the fraction of non-firm traders goes to zero ($n_{2}/n\to 0$). Since the total cost for non-firm traders is given by
@@
then taking the limit as $n_{2}/n\to 0$ yields $\frac{\kappa\lambda_{n_{2}}}{1-e^{-\kappa}}$. This means that if a firm centralizes its trades and then splits them infinitely the result is that the firm and non-firm traders in the aggregate share a total cost of $\frac{\kappa}{1-e^{-\kappa}}$ in proportion to their fractional target quantities.
@@ 6.5 Visualizations of strategic centralization implementation costs
The firm has one trader: we make three plots (each with three sub-plots) in which the firm always has one actual trader ($n_{1}=1$) and the total number of traders is 2, 3 and 10 so that the firm is 1/2, 1/3 and 1/10th of the total number of traders. We then examine implementation cost changes for three levels of alpha-decay, $\kappa=1,2.5$ and $5$. There are a few obviously interesting trends.
Figure 8: The firm has one trade ($n_{1}=1$) for relatively low alpha-decay ($\kappa=1$)for cases from left to right with non-firm traders $n_{2}=2,4,6$ and $9$. In all cases the firm’s implementation cost is lower by adding "pretending" to have more traders than it does.
Figure 9: Identical setup as Figure 8 but for $\kappa=5$. The firm has one trade ($n_{1}=1$) for relatively low alpha-decay ($\kappa=1$)for cases from left to right with non-firm traders $n_{2}=2,4,6$ and $9$. In all cases the firm’s implementation cost is lower by adding "pretending" to have more traders than it does.
Figure 10: Similar setup as Figures 8 and 9 but this time the trade is significantly more "crowded" because the total number of traders is set to much higher values. The firm still have one trader, but now the total number of traders is from left to right is 25, 50, 100 and 250. In all cases it is still the case that adding some traders lowers costs.
Firm’s traders represent most of the traders taking up the trade: Below we see the case where the firm can substantially benefit from masking its actual number of traders by consolidating its trades so as to make it appear as if there are fewer traders in the market then there actually are.
Figure 11: In this plot $\kappa=1$ so alpha-decay is relatively low. In each plot the firm has approximately 2/3rds of the total number of traders. The black dot represents the case where the firm does not centralize. One can see that consolidation (moving to the left of the black dot on the plot) lowers the cost of trading for a moderate decrease in the apparent number of traders. We see that as the total number of traders in competition grows (as seen by the different plots starting from the left at $n=15$ versus the rightmost plot where $n=50$), the available benefit from centralization decreases.
Figure 12: This is the same setup as Figure 11 except that alpha-decay is set to be significantly higher. In the prior plot, $\kappa$ was set to 1 and in this plot it is $25$. Generally speaking the pattern in the two plots is the same, however, in general the benefit from centralization is seen to be higher for this case when $\kappa$ is higher.